OpenStudy (anonymous):

I need help adding positive/negative fractions. 1/4 + -3/5?

OpenStudy (anonymous):

***subtracting

OpenStudy (anonymous):

1/4 - 3/5

OpenStudy (calculusxy):

yes subtract and take the sign of the larger absolute value

OpenStudy (calculusxy):

but 3/5 is larger than 1/4 so it's 3/5 - 1/4

OpenStudy (anonymous):

You need to get the same denominator

OpenStudy (anonymous):

Uh what am I supposed to do when it comes to be subtracting 5/20 over 12/20

OpenStudy (calculusxy):

15/20 - 5/ 20 = 10/20 = 1/2

OpenStudy (anonymous):

oh ok thank you

OpenStudy (calculusxy):

now take the sign of the larger fraction

OpenStudy (anonymous):

@calculusxy How did you get 15/20?

OpenStudy (calculusxy):

oh sorry 3 x 4 / 5 x 4 = 12/ 20

OpenStudy (anonymous):

Oh, so it comes out to be 7/20 right?

OpenStudy (calculusxy):

yes

OpenStudy (anonymous):

Ok ty

OpenStudy (calculusxy):

yw